OpenStudy (jay22):

y varies inversely with x, and x = -12 when y = 5. Find the inverse variation equation and use it to find the value of x when y = 10.

OpenStudy (ivancsc1996):

\[x=\frac{ 1 }{ y }-12-\frac{ 1 }{ 5 }\]When y=10 then x=\[-12-\frac{ 1 }{ 10 }=y=\frac{ -120-1 }{ 10 }=-11.9\]I THINK